# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ VIII Evolution Equations for Scalar Perturbations
The symbol ~$\tilde{\nabla}^{2}$ denotes the Laplace-Beltrami operator with respect to the metric $\tilde{g}_{ij}(\bm{x})$ of three-dimensional subspaces of constant time:
~$\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}:=\tilde{g}^{ij}p_{{\scriptscriptstyle(1)}\mid i\mid j},\quad g^{ij}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|i|j}=-\dfrac{\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}}{a^{2}}.$ (38)
@@
By employing the expression $\dot{p}_{\scriptscriptstyle(0)}=p_{n}\dot{n}_{\scriptscriptstyle(0)}+p_{\varepsilon}\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and eliminating the time derivatives of $\varepsilon_{\scriptscriptstyle(0)}$ and $n_{\scriptscriptstyle(0)}$ with the aid of the conservation laws (10c) and (10d), the following result is obtained:
@@
The derivation of Eqs. (37) will now be conducted. To initiate this process, the term $\dot{h}^{k}{}_{k}$ is eliminated from Eq. (12a) using the expression (19), which results in the algebraic energy density constraint equation (37a).
Given that only $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is associated with scalar perturbations, it is possible to simplify the momentum constraint equations (12b) and the momentum conservation laws (12e) by employing the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}:=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}$, (19), as an alternative to $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$. By multiplying both sides of Eqs. (12b) by $g^{ij}_{\scriptscriptstyle(0)}$ and taking the covariant derivative with respect to the index $j$, one finds
@@
By combining (41) and (42) and using (19) to eliminate $\dot{h}^{k}{}_{k}$, one obtains Eq. (37b). In consequence, the three $G^{0}_{{\scriptscriptstyle(1)}i}$ momentum constraint equations (12b) have been reformulated in the form of a first-order ordinary differential equation (37b) for the perturbed spatial Ricci scalar (18).
The momentum conservation laws (12e) are rewritten by performing the differentiation with respect to time and using (10c), then dividing by $\varepsilon_{{\scriptscriptstyle(0)}}(1+w)$. The result is as follows:
@@
Upon substituting this expression into Eq. (43), it can be seen that the momentum conservation laws can be reformulated as
@@
By taking the covariant divergence of (45) with respect to the background metric tensor (5), and employing (19) and (38), one obtains Eq. (37d).
As demonstrated in Appendix B, the dynamical equations (12c) are superfluous since the system (37) comprises the constraint equations and conservation laws, and thus its solution automatically satisfies the dynamical equations. Consequently, the derivation of the equations (37) has been successfully completed.
In Appendix A, the primary result, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}/n_{{\scriptscriptstyle(0)}}$, will be derived using the systems (10) and (37). Prior to proceeding, it is crucial to demonstrate that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the true physical energy density and particle number density perturbations, respectively. This necessitates the nonrelativistic limit being employed for the interpretation of these quantities.
IX Nonrelativistic Limit
The standard nonrelativistic limit is defined by four fundamental criteria (see, e.g., Ref. [37], Sec. 4.1):
  1. 1.
    The subspaces of constant time are flat.
  2. 2.
    When the pressure is reduced to zero, there are no local fluid flows.
  3. 3.
    Relativistic gauge transformations are reduced to Newtonian gauge transformations.
  4. 4.
    The gravitational field of a density perturbation is required to be constant.
The third requirement is of considerable importance, as it permits the selection of a coordinate system in Newtonian gravitation. This fact is not referenced in the existing literature. The fourth requirement is a direct consequence of the first and second requirement.
In the case of flat three-spaces, it is evident that ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$. In a flat flrw universe, covariant derivatives are equivalent to ordinary derivatives, thereby simplifying the expression (VII) to the following form:
@@ IX Nonrelativistic Limit
The symbol $\nabla^{2}$ represents the conventional Laplace operator. Upon substituting the expression (46) into the perturbation equations (37), and employing the relation $H:=\dot{a}/a$, the following result is obtained:
@@
In order to derive the $G^{0}_{{\scriptscriptstyle(1)}0}$ constraint equation (47a) from Eq. (37a), it is first necessary to eliminate the function $\varepsilon_{{\scriptscriptstyle(1)}}$ using the expression (4a). The result is
@@
Subsequently, the Friedmann equation (10a) with the Ricci scalar ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ set to zero, the energy conservation law (10c), and the relation between the Hubble parameter and the expansion scalar, namely, $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, are employed to obtain the result $\dot{\varepsilon}_{\scriptscriptstyle(0)}/\dot{\theta}_{\scriptscriptstyle(0)}=2H/\kappa$. Upon substituting this expression into (48), the desired result, Eq. (47a), is obtained. With the condition of ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, the first requirement is satisfied.
Subsequently, the second requirement will be implemented. Upon substituting $p=0$, i.e., $w=0$ and $\beta=0$, into the momentum conservation laws (45), the result is
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}=-2Hu^{i}_{{\scriptscriptstyle(1)}}.$ (49)
By lowering the index $i$ with the background metric $g_{{\scriptscriptstyle(0)}ij}$ given by (5), and using the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$, one finds
$\dot{u}_{{\scriptscriptstyle(1)}i}=0.$ (50)
These equations admit only functions of the spatial coordinates as solutions. As these equations are part of the system (12), it can be concluded that the gauge modes, $\psi_{|i}$, given by (28), are the only solutions of equations (50). Consequently, it can be deduced that
$p\rightarrow 0\quad\Rightarrow\quad u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0.$ (51)
In the absence of pressure, local pressure gradients are nonexistent. Consequently, the emergence of local fluid flows is precluded. This implies that
$u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0\quad\Rightarrow\quad\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})\rightarrow\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x}).$ (52)
Therefore, in the event of a zero fluid flow, the density perturbations are observed to remain static. It can be concluded that the second requirement is fulfilled by the limits (51)–(52).
Given that the physical parts of the spatial components of the fluid four-velocity are identically zero, only the gauge modes $\psi_{|i}$ remain. It can thus be concluded with confidence that the gauge modes may be taken to be equal to zero without any loss of physical information. Taking $\hat{u}_{{\scriptscriptstyle(1)}i}=\psi_{|i}=0$ implies that $\psi$ is an arbitrary infinitesimal constant. Upon substituting $\psi=C$ into the expressions (6), the result is that the relativistic gauge transformation between synchronous coordinates reduces in the limits given in (51) to the gauge transformation of Newtonian gravitation:
@@
This implies that within the Newtonian theory of gravitation, there is the possibility of shifting the time coordinate and selecting the spatial coordinates at will. In accordance with the expressions (53), the third requisite is thereby fulfilled.
The fourth and final requirement will now be discussed. Given that $u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}=0$ and $\hat{u}^{i}_{{\scriptscriptstyle(1)}}=0$, it follows that $\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k}=0$. By combining this result with (52), the linearized Einstein equations and conservation laws (47) are reduced to
$\nabla^{2}\phi(t,\bm{x})$ $=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}),$ (54a)
@@
The constraint equations (54a) and (54b) are combined to yield the following result:
$\nabla^{2}\phi(\bm{x})=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t_{0})\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}).$ (55)
The potential $\varphi(\bm{x})$ is defined as $\varphi(\bm{x}):=\phi(\bm{x})/a^{2}(t_{0})$, which results in the Poisson equation of Newtonian gravitation:
$\nabla^{2}\varphi(\bm{x})=4\pi G_{\text{N}}\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{c^{2}}.$ (56)
With (56), the fourth and final requirement has been met.
In the nonrelativistic limit, the background equations (10) are given by
@@
Consequently, the conservation laws imply that $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$. As may be observed, Eqs. (54c)–(54d) yield solely gauge modes as solutions: specifically, $\hat{\varepsilon}_{\scriptscriptstyle(1)}=C\dot{\varepsilon}_{\scriptscriptstyle(0)}$, $\hat{n}_{{\scriptscriptstyle(1)}}=C\dot{n}_{{\scriptscriptstyle(0)}}$, and $\hat{\theta}_{{\scriptscriptstyle(1)}}=C\dot{\theta}_{{\scriptscriptstyle(0)}}$, where $C$ is an infinitesimal constant of the gauge transformation (53) of Newtonian gravitation. This, combined with $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$ yields $\varepsilon_{{\scriptscriptstyle(1)}}=n_{{\scriptscriptstyle(1)}}mc^{2}$. Combining the latter two relations with the definitions (4a) one finds the well-known special relativistic relation between the rest mass density and energy density:
@@
The equations (54c) and (54d) represent the residual form of the conservation laws (47c) and (47e) in the nonrelativistic limit. It is evident that equations Eqs. (54c) and (54d) have no physical significance. These equations are decoupled from the physical constraint equations (54a) and (54b) and thus do not form part of Newtonian gravitation. Therefore, the potential $\zeta$ which appears in $\theta_{{\scriptscriptstyle(1)}}$, (24), is no longer relevant. In view of the aforementioned considerations, it becomes evident that solely the Poisson equation (56) and the special relativistic relation (58) have survived.
Following an examination of the four points regarding the nonrelativistic limit, the following conclusion is reached. It has been demonstrated that the system of linearized Einstein equations and conservation laws for scalar perturbations (37), when combined with the gauge-invariant quantities (4a), reduces in the nonrelativistic limit to the Poisson equation (56) of Newtonian gravitation and the special relativistic relation (58). This indicates that the gauge-invariant quantities $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, defined by expressions (4a), represent the local perturbations of the energy density and the particle number density, respectively, in Newtonian gravitation. Therefore, it can be stated that $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ represent the actual local density perturbations of the energy density and the particle number density in general relativity. In conclusion, since the gauge-invariant perturbation of the expansion, $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ (4b), is identically zero, it can be stated that density perturbations have no influence on the expansion of the universe.
The subsequent section of this study will proceed to derive the evolutionary equations for $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$. These equations permit the investigation of the evolution of density fluctuations in closed, flat, and open flrw universes independently of the choice of coordinate system.
X Evolution Equations for Density Fluctuations

### v3

@@ IX Evolution Equations for Scalar Perturbations
[…]
The symbol $\tilde{\nabla}^{2}$ denotes the Laplace-Beltrami operator with respect to the metric $\tilde{g}_{ij}(\bm{x})$ of three-dimensional subspaces of constant time:
$\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}:=\tilde{g}^{ij}p_{{\scriptscriptstyle(1)}\mid i\mid j},\quad g^{ij}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|i|j}=-\dfrac{\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}}{a^{2}}.$ (38)
@@
Using the expression $\dot{p}_{\scriptscriptstyle(0)}=p_{n}\dot{n}_{\scriptscriptstyle(0)}+p_{\varepsilon}\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and eliminating the time derivatives of $\varepsilon_{\scriptscriptstyle(0)}$ and $n_{\scriptscriptstyle(0)}$ with the aid of the conservation laws (11c) and (11d), the following result is obtained:
@@
We will now derive Eqs. (37). To begin, the term $\dot{h}^{k}{}_{k}$ is eliminated from Eq. (13a) using the expression (20). This results in the algebraic energy density constraint equation (37a).
Given that only $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is associated with scalar perturbations, it is possible to simplify the momentum constraint equation (13b) and the momentum conservation law (13e) by using the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}:=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}$}, (20), as an alternative to $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$. By multiplying both sides of Eq. (13b) by $g^{ij}_{\scriptscriptstyle(0)}$ and taking the covariant derivative with respect to the index $j$, one finds
@@
By combining (41) and (42) and using (20) to eliminate $\dot{h}^{k}{}_{k}$, one obtains Eq. (37b). Consequently, the $G^{0}_{{\scriptscriptstyle(1)}i}$ momentum constraint equation (13b) has been reformulated as a first-order ordinary differential equation (37b) for the perturbed spatial Ricci scalar (VII.1) expressed in the potentials.
The momentum conservation law (13e) is reformulated by first performing the differentiation with respect to time. Then, we use Eq. (11c) to eliminate $\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$. Next, dividing the result by $\varepsilon_{{\scriptscriptstyle(0)}}(1+w)$ yields
@@
Upon substituting this expression into Eq. (43), the momentum conservation law becomes
@@
Taking the covariant divergence of (45) with respect to the spatial background metric tensor (8), and using (20) and (38) yields Eq. (37d).
As demonstrated in Appendix B, the dynamical equation (13c) is redundant since the system (37) comprises the constraint equations and conservation laws. Thus, the derivation of Eqs. (37) has been completed.
X Newtonian Limit
As demonstrated in Sec. IV, the gauge-invariant quantities, $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, which are intended to describe density perturbations, can only be defined in one way. This section proves that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent perturbations to energy and particle number densities, respectively.
In Newtonian gravity, space is Euclidean. Consequently, the Newtonian limit can only be realized in a flat flrw universe. In this case, the Newtonian limit is reached when gravity is weak and the velocities of the particles are negligible compared to the speed of light. These conditions are met in a linear perturbation theory and when pressure approaches zero.
In a flat flrw universe where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, covariant derivatives are equal to ordinary derivatives. This simplifies the expression (VII.1) to
@@ X Newtonian Limit
The symbol $\nabla^{2}$ represents the conventional Laplace operator. By substituting the expression (46) into the perturbation equations (37) and using that $H:=\dot{a}/a$, we obtain the result:
@@
To derive the $G^{0}_{{\scriptscriptstyle(1)}0}$ constraint equation (47a) from Eq. (37a), the function $\varepsilon_{{\scriptscriptstyle(1)}}$ must first be eliminated using the expression (6a). The result is
@@
Setting ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$ in the Friedmann equation (11a) and differentiating the result with respect to time, and using the relation $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, we obtain $\dot{\varepsilon}_{\scriptscriptstyle(0)}/\dot{\theta}_{\scriptscriptstyle(0)}=2H/\kappa$. Substituting this expression into Eq. (48) yields the desired result, Eq. (47a). The Newtonian limit can now be derived using the system (47).
In a universe filled with a nonrelativistic fluid, we have $w\ll 1$ and $\beta^{2}\ll 1$. Neglecting the quantities $w$ and $\beta^{2}$ with respect to the constants of order one in the momentum conservation law (45), yields
The second equation is derived by lowering the index $i$ in the first equation using the background metric $g_{{\scriptscriptstyle(0)}ij}$ and the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$. Accordingly, the evolution of density perturbations is highly dependent upon the presence of pressure gradients, represented by $p_{{\scriptscriptstyle(1)}|j}$, where $p_{{\scriptscriptstyle(1)}}$ is given by (14). In a pressureless fluid, i.e., when $w=0$ and $\beta=0$, however, the result is
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}+2Hu^{i}_{{\scriptscriptstyle(1)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}=0.$ (50)
The second equation has solutions that are only functions of the spatial coordinates. Since these equations are part of the system (13) that is invariant under the gauge transformation (3) with $\xi^{\mu}(t,\bm{x})$ given by (9), it can be inferred that the gauge modes given by (29), are the only solutions to equations (50). Therefore, it can be concluded that
$p\rightarrow 0\quad\Rightarrow\quad u_{{\scriptscriptstyle(1)}i}^{{\textnormal{phys}}}(t,\bm{x})\rightarrow 0.$ (51)
In other words, in the absence of pressure, local pressure gradients $p_{{\scriptscriptstyle(1)}|i}$ are nonexistent, so local fluid flows are impossible. Thus, in the Newtonian limit, density perturbations are independent of time, that is, they are static. The remaining part of this section examines the consequences of the limit (51).
Since the components (51) of the fluid velocity are zero, only the gauge modes $\hat{u}_{{\scriptscriptstyle(1)}i}=\psi_{|i}$ remain. We can set these gauge modes to zero without losing any physical information. Substituting $\psi_{|i}=0$ into the expressions (9) yields $\xi^{0}=C$ and $\xi^{i}=\chi^{i}$, implying that the relativistic gauge transformation (3) reduces to the gauge transformation of Newtonian gravity:
@@
Thus, we arrive at the well-known fact that in Newtonian gravity, we can freely select the spatial coordinates and shift the time coordinate at will, and that space and time are independent of each other.
Since both $\hat{u}_{{\scriptscriptstyle(1)}i}=0$ and $u_{{\scriptscriptstyle(1)}i}^{{\textnormal{phys}}}=0$ we find from (20) that $\vartheta_{\scriptscriptstyle(1)}=0$. By substituting the latter into the system (47) we obtain
$\nabla^{2}\phi(t,\bm{x})$ $=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x}),$ (53a)
@@
The constraint equations (53a) and (53b) imply
$a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=a^{2}(t_{0})\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t_{0},\bm{x}).$ (54)
This equation applies at any given time $t_{0}$. From the Friedmann equation (11a), it follows that, for ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, one may take $a(t_{0})=1$, so that
$a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}).$ (55)
This expression is consistent with the limit (51) that results in static density perturbations. Combining (53a), (53b), and (55) yields
$\nabla^{2}\phi(\bm{x})=4\pi G_{\text{N}}\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{c^{2}},$ (56)
which is the Poisson equation of Newtonian gravity.
Finally, the meaning of $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is established. For a universe filled with a pressureless fluid, the background equations (11) are given by
@@
Consequently, the conservation laws imply that $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$. As can be verified by substitution, Eqs. (53c)–(53d) yield solely gauge modes as solutions: specifically, $\hat{\varepsilon}_{\scriptscriptstyle(1)}=C\dot{\varepsilon}_{\scriptscriptstyle(0)}$, $\hat{n}_{{\scriptscriptstyle(1)}}=C\dot{n}_{{\scriptscriptstyle(0)}}$, and $\hat{\theta}_{{\scriptscriptstyle(1)}}=C\dot{\theta}_{{\scriptscriptstyle(0)}}$, where $C$ is an infinitesimal constant of the gauge transformation (52) of Newtonian gravity. This, combined with $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$ yields $\varepsilon_{{\scriptscriptstyle(1)}}=n_{{\scriptscriptstyle(1)}}mc^{2}$. Combining the latter two relations with the definitions (6a) one finds
@@
This is the well-known special relativistic relation between rest mass and energy.
@@
The equations (53c) and (53d) represent the residual form of the conservation laws (47c) and (47e) in the Newtonian limit. Since Eqs. (53c) and (53d) only have gauge modes as solutions, they clearly have no physical significance. These equations are decoupled from the physical constraint equations (53a) and (53b) and thus do not form part of Newtonian gravity. Therefore, the potential $\zeta$ which appears in $\theta_{{\scriptscriptstyle(1)}}$ (25) is no longer relevant. This is consistent with the comment just below (25). Thus, in the Newtonian limit, only the Poisson equation (56) and the special relativistic relation (58) remain.
We now come to a final conclusion. The analysis above demonstrates that, in the Newtonian limit, only the Poisson equation (56) and the relation (58) remain. Therefore, $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ (6a) represent the perturbations in energy and particle number density, respectively, in Newtonian gravity. Consequently, they also represent perturbations in energy and particle number density in general relativity. Finally, we note that, since $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$ (6b), density perturbations do not affect the expansion of the universe.
In the next subsections, relative density perturbations in the Newtonian limit are discussed, and the perturbation theory based on Newtonian gravity is examined in the context of relativity theory.
X.1 Relative Density Perturbation in the Newtonian Limit
The expression (55) implies that
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{a^{2}(t)}.$ (59)
Since Eq. (11c) with $w=0$ implies $\varepsilon_{{\scriptscriptstyle(0)}}\propto a^{-3}$, it follows that the relative density perturbation $\delta_{\varepsilon}$ evolves according to
$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{{\scriptscriptstyle(0)}}(t)}\propto\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})a(t).$ (60)
Thus, due to the expansion of the universe, the relative density perturbation $\delta_{\varepsilon}$ for a static density perturbation $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})$ increases proportionally to the scale factor $a(t)$ of the universe. In other words, a static density perturbation becomes relatively denser when viewed against the decreasing energy density $\varepsilon_{{\scriptscriptstyle(0)}}$ in the background universe.
X.2 Newtonian Gravity
As demonstrated in Appendix C, the relative perturbation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ is gauge-dependent. The gauge mode associated with $\delta$ is given by (152a). For $w\ll 1$, it is a solution to the homogeneous part of Eq. (157a). This gauge mode does not disappear from the solution because $H\neq 0$ in a nonempty universe. Consequently, the quantity $\delta$ depends on the choice of coordinates and therefore has no physical meaning. As can be seen from Eqs. (57), $H\neq 0$ even in the Newtonian limit. Therefore, $\delta$ must also depend on the choice of coordinates in that limit and consequently has no physical significance. This is consistent with the fact that, in the Newtonian limit, the relativistic gauge transformation (3) converges to the Newtonian gauge transformation (52). Given that the system (151) reduces to the system (157) for a nonrelativistic fluid, it can be concluded that the system (157) is inadequate for examining the evolution of density perturbations.
The existing literature derives Eq. (156), from Newtonian gravity under two conditions: the expansion of the universe must be considered and the equation of state must be nonrelativistic, i.e., $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}\ll 1$ and $\beta^{2}\ll 1$. See, for example, Ref. [7], Sec. 6.2, Ref. [42], Sec. 15.9, and Ref. [48], Eq. (3.17). However, the gauge transformation (52) shows that $\delta$ depends on the choice of coordinates in Newtonian gravity. Therefore, it also lacks physical significance in Newtonian gravity. Furthermore, a comparison of the conventional equation (156) and the exact equation (157a) reveals that equation (156) is incomplete. Consequently, the conventional evolution equation (156) — whether derived from general relativity or Newtonian gravity modified to account for the expansion of the universe — also has gauge-dependent solutions. These findings indicate that Newtonian gravity is inadequate in accurately describing the evolution of cosmological density perturbations. Therefore, there exists no distinction between sub- and super-horizon density perturbations.
XI Relativistic Cosmological Perturbation Theory